You roll a pair of dice, one is red and the other is green. What is the probability of rolling a 5 on the red die and an even number on the green one?
Answer by stanbon(75887)

You roll a pair of dice, one is red and the other is green. What is the probability of rolling a 5 on the red die and an even number on the green one?
------
What happens on the red and what happens on the green are
independent events.
P(red 5 AND green even) = (1/6)(3/6) = 3*36 = 1/12
=============================
Cheers,
Stan H.
============